OpenStudy (anonymous):

Rewrite the equation in a rotated x'y'-system without an x'y' term. Express the equation involving x' and y' in the standard form of a conic section. 31x2 + 10xy + 21y2 -144 = 0

OpenStudy (anonymous):

angle = 30 so x = (sqrt3x - y)/2 y = (x + sqrt3y)/2

OpenStudy (anonymous):

so i plg in x and y and i now have this long really confusing equation that im having trouble simplifying..... did i even take the right steps?

OpenStudy (anonymous):

Let me see if I can remember how to do... We have Ax^2 + Bxy + Cy^2 +F = 0 whre A = 31 B = 10 C = 21 D = -144 A' = A cos^2 theta + B sin theta cos theta + C sin^s theta C' = A sin^2 theta - B sin theta cos theta + C cos^2 theta F' = F and where Tan 2 theta = (B/(A-C))

OpenStudy (anonymous):

B = 10sqrt 3 cot 2 theta = ((A - C)/B) cot 2 theta = 1/sqrt 3 = 60 degrees/ 2 = 30 degrees

OpenStudy (anonymous):

atleast that's the way ive learned it so far. and then x = x'cos theta - y'sin theta y = x'sin theta + y'cos theta

OpenStudy (anonymous):

cot 2 theta = ((A - C)/B) cot 2 theta = 1/sqrt 3 = 60 degrees/ 2 = 30 degrees This is the same thing as mine.......

OpenStudy (anonymous):

oh ok. but then what's next do i plug in those x and y values?

OpenStudy (anonymous):

You need to do this (in my way) A' = A cos^2 theta + B sin theta cos theta + C sin^s theta C' = A sin^2 theta - B sin theta cos theta + C cos^2 theta F' = F and you can get all the trigonometry from a triangle given by Tan 2 theta = (B/(A-C)) = 1 Where do you get 30 degrees?

OpenStudy (anonymous):

A-C/ B is 1 right?

OpenStudy (anonymous):

i got 1/ sqrt3

OpenStudy (anonymous):

31-21/10 ?

OpenStudy (anonymous):

i included the sqrt3 so that B = 10(sqrt3)

OpenStudy (anonymous):

What sqrt 3?

OpenStudy (anonymous):

omg! i must have written that down on my paper different. you're right there's no sqrt3 that changes everything! haha

OpenStudy (anonymous):



OpenStudy (anonymous):

OK, let me know if it still doesn't work out...:-)

OpenStudy (anonymous):

ok thank you